OpenStudy (anonymous):

A contractor digs a hole for the basement of a new house that is 40 feet long and 10 1/2 feet deep. What is the width of the hole if 6,510 cubic feet of dirt are removed?

OpenStudy (anonymous):

15.5

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

to find the volume of an object, you must multiply the 3d figure's dimensions which is depth, width, length. since the length, depth, and volume are found you must find the width using the expression of [L*D]/V=W

OpenStudy (anonymous):

im not getting it 40x10.5=42000/6510=6.45

OpenStudy (jdoe0001):

\(\begin{array}{llll} length\cdot &width\cdot &height=&volume\\ \hline\\ 40&\bf w&10.5=&6510 \end{array}\) solve for "w"

OpenStudy (anonymous):

i got 6.45

OpenStudy (anonymous):

the choice has no 6.45

OpenStudy (jdoe0001):

6.45?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

EXACTLY

OpenStudy (jdoe0001):

hmmm from lwh=v what would "w" be?

OpenStudy (anonymous):

W=15.5

OpenStudy (anonymous):

how is 6.45=15.5???????????

OpenStudy (jdoe0001):

is not, because 40 * w * 10.5 = 6510 \(\ne 6.45\)

OpenStudy (anonymous):

40*10.5=420 6510/420=15.5

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i divided the wrong way

OpenStudy (anonymous):

do u mind giving me another question?

OpenStudy (anonymous):

SURE

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

TY ?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

need a question

OpenStudy (anonymous):

IT IS ON THE LIST 3RD

OpenStudy (anonymous):

what do you mean (your on CAPS LOCK)

OpenStudy (anonymous):

here is the question

OpenStudy (anonymous):

the highest temprature recorded in the town of westgate this summer was 0 Fahrenheit. last winter, the lowest temprature recorded was -1 fahrenheit. find the difference between these extremes.